# StereoDiffusion: Training-Free Stereo Image Generation Using Latent Diffusion Models

### v1

@@ 4.3 Ablation study
5 Limitations and Discussion
Our method still relies on the disparity map. If the results generated by other depth estimation models are inaccurate, our method will also be unable to produce high-quality stereo images. Furthermore, when using high-precision disparity maps obtained from actual device measurements, the results may not be entirely satisfactory, as shown in Table 3 and Fig. 7.
When using our depth to stereo image model, one may observe overlapping areas in the generated images. This issue might stem from the LatentDepth2ImageDiffusion model we used, which tends to fill blank areas with pixels from adjacent main subjects rather than background elements. In such cases, a better-quality image can be generated by first generating a single image using the Depth2Image model, and then applying our Image to Stereo Image Pairs method, as illustrated in Fig. 8.
We found that our method can be used for inpainting tasks with the original text prompt to an image Stable Diffusion model. We conducted a simple test where, after obtaining $\bm{x}_{t}$ using null-text inversion, we applied various masking ratios to the right side and tested whether Stable Diffusion could fill in the blank areas within the mask during denoising. The results, as shown in Fig. 9, indicate that our method is somewhat effective for inpainting when a smaller area of the image is masked. However, when a larger portion of the image is masked, the inpainting results exhibit a strong patchwork appearance. Applying our method to inpainting tasks might require further modifications to both the model and the technique.
6 Conclusion
Appendix A Quantitative evaluation experiments setting
After rendering with these settings, we obtained the left and right images of the stereo image pairs.
Given that Stable Diffusion can only generate images of $512\times 512$ resolution, and the Middlebury dataset images are about 5 million pixels, we scaled both the dataset images and the corresponding depth maps to $512\times 512$. For the Middlebury dataset, whose groundtruth disparity maps are noisy, we applied a Gaussian blur with a radius of 3 to smooth the disparity maps. Regarding the KITTI dataset, where the image size is $375\times 1242$ with an aspect ratio of approximately 3.3, directly scaling images to $512\times 512$ could lead to excessive stretching, negatively impacting many models’ performance. Therefore, we proportionally scaled the images to $512\times 1696$ and then applied a center crop to $512\times 512$ to avoid excessive stretching. As null-text inversion technique is required, we used the Stable Diffusion version 1.5 for this test, setting the denoising steps to 50.
For the 3D photography method [12], we used the disparity map generated by the integrated MiDaS model [15] within its framework instead of the groundtruth disparity map. This was due to the extensive time required—up to two hours—for mesh reconstruction of a single image using the groundtruth disparity map with 3D photography. We hypothesize that this inefficiency arises when 3D photography attempts to reconstruct stereo image pairs from the disparity map, necessitating operations like breaking up discontinuous vertices in the mesh. Such processes become computationally intensive when the groundtruth disparity map is excessively noisy, leading to a proliferation of isolated vertices that consume substantial CPU resources. For the purpose of benchmarking and considering the rarity of obtaining groundtruth disparity maps in practical scenarios, we evaluated the results using both groundtruth disparity maps (denoted as GT disparity) and pseudo disparity maps generated by depth estimation models (denoted as Pseudo disparity). The depth estimation model we employed was DPT [14]. Since the use of Deblur results in lower scores, neither method employed deblur; details can be found in Section 4.3 Ablation study. When creating stereo image pairs using RePaint [45], we generate a mask for the blank areas left after moving the left-side image and then perform inpainting on the masked areas. The model inet256 we utilized for this purpose was trained on ImageNet. Since RePaint’s maximum supported output image size is $256\times 256$, we downsized the images to $256\times 256$ before conducting inpainting. However, considering that all other methods are evaluated at a $512\times 512$ resolution, for fairness, we only upscale the inpainted area within the mask from $256\times 256$ to $512\times 512$, while maintaining the original resolution for the area outside the mask.
@@ Appendix A Quantitative evaluation experiments setting
Appendix B Attempts of fine-tuning Stable Diffusion model to genreate stereo image pairs
@@
Figure 10: Example of images generated by stereo fine-tuned Stable Diffusion: The images reveals that while the generated left and right images exhibit certain similarities, the extent of this resemblance falls significantly short of the requirements for stereo imaging. Even during training, maintaining pixel-level consistency between the left and right images proves challenging, and the quality of images generated during test exhibits notable deficiencies.
@@
ControlNet [35], known for its capability to manipulate the posture of images generated by Stable Diffusion, produces images that are structurally similar to the input image but with different content. We hypothesized that this might be beneficial for generating stereo images. Consequently, we adopted an architecture similar with ControlNet. A neural network block $F(\cdot;\Theta)$ with a set of parameters $\Theta$ transforms a feature map $\bm{x}$ into another feature map $\bm{y}$.
$\bm{y}=F(\bm{x};\Theta)$ (13)
We have frozen all the parameters $\Theta$ of the original Stable Diffusion and created a trainable copy $\Theta_{c}$. The neural network blocks are interconnected through a distinctive convolution layer, which is initialized with zero weights and biases. The operation can be represented by the following equation
$\bm{y}_{\mathrm{c}}=\mathcal{F}(\bm{x};\Theta)+\mathcal{Z}\left(\mathcal{F}\left(\bm{x}+\mathcal{Z}\left(\bm{c};\Theta_{\mathrm{z}1}\right);\Theta_{\mathrm{c}}\right);\Theta_{\mathrm{z}2}\right)$ (14)
@@
Using ControlNet only maintains the general content of the images, which is insufficient for generating stereo image pairs. We aim for Stable Diffusion to generate stereo image pairs concurrently. To achieve this, we align even-numbered images in the batch with their adjacent odd-numbered counterparts, such as 0 with 1, and 1 with 2, to create a stereo effect between each adjacent pair. Inspired by VideoLDM[23], we introduce a 3D convolution layer and a temporal attention layer into the Stable Diffusion architecture. These layers are added after Stable Diffusion’s existing spatial layers in the U-Net. The function of 3D convolution layer’s is to break the information isolation between different samples in the same batch. Before feeding the intermediate features to the 3D convolution layer, we reshape the features from [b c h w] to [b/2 2 c h w], where b, c, h, w represent batch size, color channel, height, and width, respectively. The 2 in the reshaped second item represents the left and right images, allowing the newly added 3D convolution block to learn the distribution of the left and right stereo image pairs. The structure of the temporal attention layer is same as that in Stable Diffusion, assisting the 3D convolution layer in distinguishing different timesteps during the denoise process.
@@
Appendix C Ablation of Bidirectional attention and Stereo Pixel Shift
@@
Incorporating Bidirectional Attention and applying Stereo Pixel Shift to the both left and right latent variables can alter the original image, making it unsuitable for quantitative analysis. Therefore, we only partially showcase the results of the text prompt to stereo image generation, as depicted in Fig. 11. The simultaneous application of Bidirectional Attention and Stereo Pixel Shift to both left and right latent variables may induce changes in the original image. These modifications are currently uncontrollable. However, this may suggest a new potential of our approach: a method of controlling the generated images, akin to ControlNet, but without the need for fine-tuning.
Appendix D User test images
Figure 12: Comparison of stereo image generation techniques. RePaint et al. [45] indicates using their inpainting model to fill the blank area. HINT: The images can be viewed using the autostereogram technique to achieve a 3D effect. (Keep your eyes steady and maintain the unfocused gaze, try adjusting eyes’ focus and the distance between the autostereogram and your eyes slightly.)
In Fig. 12 we show the example images used for our user evaluation.

### v2

@@ 9 Analysis of Ablation
10 Attention module modification details
In the case of self-attention, non-rigid editing cannot be performed as the semantic layout and structures are maintained. Similar to sharing semantic information between different samples in the same batch using 3D convolution to align content across batches in video generation tasks [37, 2, 7], applying self-attention between samples within the same batch has a comparable effect [42, 4]. Querying the left-side image using the key and value of the right-side image in a unidirectional manner, enhancing the alignment from right image to left, is termed unidirectional self-attention. In contrast, employing queries from both the left and right sides to mutually query each other is referred to as bidirectional self-attention. However, bidirectional self-attention has a significant drawback: it aligns the left and right images with each other, thereby altering the input left-side image. Although this can enhance alignment, it is not a suitable option when users wish to keep the input image unchanged. Thus, despite its potential to improve alignment, the bidirectional approach may not be preferable if it is crucial to maintain the integrity of the input image. The algorithm is explained in the appendix.
@@ 10 Attention module modification details
11 Attempts of fine-tuning the Stable Diffusion model to generate stereo image pairs
@@
Figure 10: Example of images generated by stereo fine-tuned Stable Diffusion: The images reveals that while the generated left and right images exhibit certain similarities, the extent of this resemblance falls significantly short of the requirements for stereo imaging. Even during training, maintaining pixel-level consistency between the left and right images proves challenging, and the quality of images generated during tests exhibits notable deficiencies.
@@
ControlNet [4], known for its capability to manipulate the posture of images generated by Stable Diffusion, produces images that are structurally similar to the input image but with different content. We hypothesized that this might be beneficial for generating stereo images. Consequently, we adopted an architecture similar to ControlNet. A neural network block $F(\cdot;\Theta)$ with a set of parameters $\Theta$ transforms a feature map $\bm{x}$ into another feature map $\bm{y}$.
$\bm{y}=F(\bm{x};\Theta)\,.$ (12)
We have frozen all the parameters $\Theta$ of the original Stable Diffusion model and created a trainable copy $\Theta_{c}$. The neural network blocks are interconnected through a distinctive convolution layer, which is initialized with zero weights and biases. The operation can be represented by the following equation
$\bm{y}_{\mathrm{c}}=\mathcal{F}(\bm{x};\Theta)+\mathcal{Z}\left(\mathcal{F}\left(\bm{x}+\mathcal{Z}\left(\bm{c};\Theta_{\mathrm{z}1}\right);\Theta_{\mathrm{c}}\right);\Theta_{\mathrm{z}2}\right),$ (13)
@@
Using ControlNet only maintains the general content of the images, which is insufficient for generating stereo image pairs. We aim for Stable Diffusion to generate stereo image pairs concurrently. To achieve this, we align even-numbered images in the batch with their adjacent odd-numbered counterparts, such as 0 with 1, and 1 with 2, to create a stereo effect between each adjacent pair. Inspired by VideoLDM[2], we introduce a 3D convolution layer and a temporal attention layer into the Stable Diffusion architecture. These layers are added after Stable Diffusion’s existing spatial layers in the U-Net. The function of 3D convolution layers is to break the information isolation between different samples in the same batch. Before feeding the intermediate features to the 3D convolution layer, we reshape the features from [b c h w] to [b/2 2 c h w], where b, c, h, w represent batch size, color channel, height, and width, respectively. The 2 in the reshaped second item represents the left and right images, allowing the newly added 3D convolution block to learn the distribution of the left and right stereo image pairs. The structure of the temporal attention layer is same as that in Stable Diffusion, assisting the 3D convolution layer in distinguishing different timesteps during the denoise process.
@@
12 Ablation of Bidirectional attention and Stereo Pixel Shift
@@
Incorporating Bidirectional Attention and applying Stereo Pixel Shift to both the left and right latent variable can alter the original image, making it unsuitable for quantitative analysis. Therefore, we only partially showcase the results of the text prompt to stereo image generation, as depicted in Fig. 11. The simultaneous application of Bidirectional Attention and Stereo Pixel Shift to both left and right latent variable may induce changes in the original image. These modifications are currently uncontrollable. However, this may suggest a new potential of our approach: a method of controlling the generated images, akin to ControlNet, but without the need for fine-tuning.
13 User test images
Figure 12: Comparison of stereo image generation techniques. RePaint et al. [16] indicates using their inpainting model to fill the blank area. HINT: The images can be viewed using the autostereogram technique to achieve a 3D effect. (Keep your eyes steady and maintain an unfocused gaze, try adjusting the eyes’ focus and the distance between the autostereogram and your eyes slightly.)
In Fig. 12, we show the example images used for our user evaluation.